OpenStudy (anonymous):

QUADRATIC EQUATION HELP! -Z^2=-2Z-35 I know how to put them in the formulat x=-b+-[b2-4ac/2a the problem is after i do about 2 step into it i get completely lost. (stupid variable) z confuses me. please help!

OpenStudy (anonymous):

\[-z^2=-2z-35\]

OpenStudy (anonymous):

quadratic formula

OpenStudy (amistre64):

change z to x; the variable really doesnt care what it looks like :)

OpenStudy (amistre64):

\[-x^2=-2x-35\] now put it into the proper format; get it all to one side \[0=x^2-2x-35\] now you can easily see your abc stuff

OpenStudy (anonymous):

ok. so this is what i have x=-(-2)+-\[_(-2)-+\sqrt{(-2)^{2}-4z(140)}/2x\]

OpenStudy (anonymous):

\[x=2+-\sqrt{9-4x(140}/2x\]

OpenStudy (anonymous):

wait i have something wrong. its -4 not 9

OpenStudy (amistre64):

there are no "variables" in the quad formula, they all get filled in by the proper coefficients \[\frac{(-2)\pm\sqrt{(-2)^{2}-4(1)(35)}}{2(1)}\]

OpenStudy (anonymous):

oooohhhhhhh ok! now i get it! i know how to do the rest! thank you!

OpenStudy (amistre64):

:) good luck

OpenStudy (anonymous):

one question though

OpenStudy (anonymous):

wouldnt the 35 be negative?

OpenStudy (amistre64):

it would, and i also missed the negative sign on the "b" in front. good eye

OpenStudy (anonymous):

yes. either way. i know what i amo doing now. =) i appreciate the help! thanks a lot!

OpenStudy (amistre64):

yw